Cobra Golf Co.is considering a proposal to replace an existing casting machine for producing a new line of low quality golf clubs. The machine is expected to have a four-year useful life and will be depreciated according to 3-year MACRS (.25, .38, .37). The machine will cost the company $100,000 plus freight and installation costs of $20,000. The machine will be fully depreciated and will have an ending market value of $30,000. Expanding the product line will increase inventories by $10,000, but costs will decrease by $50,000 per year. Assume a tax rate of 40 percent and a cost of capital of 10 percent. What are the annual cash flows? And calculate the NPV and IRR.
Answer #1

Answer.

1. Calculation of annual cash flow

= Decrease in cost per year - Increase in invetories cost per year

= $50000 - $10000

  Annual cash flow = $40,000

2. Calculation of NPV

Working note 1

Total cost of asset = Cost of purchae+Freight &Installation charges

= $100000+20000

Total cost of asset =$1,20,000

Woking note 2-

Calculation of tax shield on depreciation

Year

Cost of

asset

(a)

Depreciation

rate as per MACRs

(b)

Depreciation

amount

(c=a*b)

Tax shield

@ 40%

(d=c*40%)

1 $120000 25% $30000 $12000
2 $120000 38% $45600 $18240
3 $120000 37% $44400 $17760

Calculation of NPV

Year Particulars

Cash Flow

(a)

Discountin

Factor @ 10%

(b)

Present Value

(c= a*b)

0 Cost of asset -$1,20,000 1.00 -$1,20,000
1-4 Cash flow (wn.1) $40,000 3.17 $1,26,800
4

Salvage value

($30000(1-0.40)

(Refer note 1 below)

$18000 0.683 $12,294
1

Tax shield on depreciation

(wn.2)

$12000 0.909 $10,908
2 Tax shield on depreciation $18240 0.826 $15,066.24
3 Tax shield on depreciation $17760 0.751 $13,337.76
NPV $58,406

Note 1- It is assumed that salvage value of asset is subject to tax deduction.

3. Calculation of IRR

at IRR,

Total Cash inflow = Total cash outflow

Total cash outflow = $1,20,000

Total cash inflow = Net cash flow of 4 years + Salvage value

= ($40000*4)+$18000

Total cash inflow = $1,78,000

Avg. cash flow = $1,78,000/4 years

= $44,500

Present value annity factor(PVAF) for IRR = Initial invetsment /Avg. cash flow

PVAF = $1,20,000 / $44,500

     PVAF = $ 2.70 (approx.)

Hence,

PVAF 16% for 1-4 year = 2.80

PVAF 18% for 1-4 year = 2.69

Hence IRR is between 16%-18%

Calculation of IRR

Year Particulars

Cash Flow

(a)

Discountin

Factor @16%

(b)

Present

Value

(c=a*b)

Discounting

Factor @ 18%

(d)

Present

Value

((e=a*d)

0 Cost of asset -$1,20,000 1.00 -$1,20,000 1.00 -$1,20,000
1-4 Cash flow $40,000 2.80 $,1,12,000 2.69 $1,07,600
4 Salvage Value $18,000 0.552 $9,936 0.516 $,9288
NPV $1936 -$3112

IRR,

=Start Rate+[surplus /surplus-(-defeciet)]*(end rate-start rate)

=16+[$1936/ $1936-(-$3112)] *(18-16)

= 16+ [$1936/$5048]*2

= 16+0.38*2*100

IRR = 16.76%

Note - Depreciation in not considered while calculating IRR because it is non cash item